Non-Constant Growth Valuation: Divining Rod Water Resources Corp (DRWR) has experienced rapid growth that is expected...

  1. Non-Constant Growth Valuation: Divining Rod Water Resources Corp (DRWR) has experienced rapid growth that is expected to continue for the next several years. Specifically, high growth of 30% for the next three years, followed by constant growth of 6% thereafter is expected. The most recent dividend was $2.50 and the required rate of return is 14%. What is the intrinsic value of DRWR’s stock?
Answer #1

D1=(2.5*1.3)=3.25

D2=(3.25*1.3)=4.225

D3=(4.225*1.3)=5.4925

Value after year 3=(D3*Growth rate)/(Required return-Growth rate)

=(5.4925*1.06)/(0.14-0.06)

=72.775625

Hence intrinsic value=Future dividend and value*Present value of discounting factor(rate%,time period)

=3.25/1.14+4.225/1.14^2+5.4925/1.14^3+72.775625/1.14^3

=$58.93(Approx).